A $15,000 investment earned $375 in interest in 3 months. Find the annual simple interest rate on the loan
If you travel 17 mi east and then 11 mi north, how far are you from your starting point? Round to the nearest tenth.

Answer by rfer(16322)

375=15000*r*0.25
375/3750=r
0.10=r
r=10%
-----------------
17^2+11^2=c^2
289+121=c^2
410=c^2
sqrt 410=c
c=20.2 mi